Given the demand function PD=27-Q²and supply function PS=2Q+3 Assuming perfect competition find the consumer surplus and producer surplus
1
Expert's answer
2020-10-02T06:22:05-0400

At equilibrium

"PD=PS"

"(27-Q^2)=(2Q+3)"

take like terms on same side

"27-3=2Q+Q^2"

"24=2Q+Q^2"

"Q^2+2Q-24=0"

Write 2Q as difference.

"Q^2+6Q-4Q-24=0"

Factor Q and 4 out

"Q(Q+6)-4(Q+6)"

Factor (Q+6) out of expression

"(Q+6)(Q-4)"

Hence

"Q+6=0"

"Q-4=0"

Hence Q = 4 or -6

(Note point of equilibrium should be positive)


Q at equilibrium = 4


P at equilibrium

[27-42] = [2(4)+3] = 11


The consumer surplus is,

"\int(27-Q^2)dQ - (11)(4)"

"[27Q - Q^3\/3]- 44"

Since Q = 4

"[27(4)-4^3\/3]-44" = 42.67=$42.67


The producer surplus is

"(11)(4)-\int(2Q+3)dQ"

"44-[2Q^2\/2+3Q)"

Since Q=4

"44-[2(4^2)\/2+12]= 16" =$16